# Surgical-LVLM: Learning to Adapt Large Vision-Language Model for Grounded Visual Question Answering in Robotic Surgery

### v2

@@ Abstract
Recent advancements in Surgical Visual Question Answering (Surgical-VQA) and related region grounding have shown great promise for robotic and medical applications, addressing the critical need for automated methods in personalized surgical mentorship. However, existing models primarily provide simple structured answers and struggle with complex scenarios due to their limited capability in recognizing long-range dependencies and aligning multimodal information. In this paper, we introduce Surgical-LVLM, a novel personalized large vision-language model tailored for complex surgical scenarios. Leveraging the pre-trained large vision-language model and specialized Visual Perception LoRA (VP-LoRA) blocks, our model excels in understanding complex visual-language tasks within surgical contexts. In addressing the visual grounding task, we propose the Token-Interaction (TIT) module, which strengthens the interaction between the grounding module and the language responses of the Large Visual Language Model (LVLM) after projecting them into the latent space. We demonstrate the effectiveness of Surgical-LVLM on several benchmarks, including EndoVis-17-VQLA, EndoVis-18-VQLA, and a newly introduced EndoVis Conversations dataset, which sets new performance standards. Our work contributes to advancing the field of automated surgical mentorship by providing a context-aware solution. Our code is available at github.com/gkw0010/Surgical-LVLM.
I Introduction
@@
The progress achieved in surgical vision question answering (VQA) and related region grounding has attracted many researchers since these techniques demonstrate great potential for applications in robotics and healthcare. To integrate the above techniques and achieve them simultaneously, [1] pioneered the surgical visual question localized-answering (VQLA), which predict answers with localized results through the input text and images. Current approaches in this field predominantly rely on the expertise of expert surgeons, often failing to acknowledge the challenges posed by real-world constraints, such as the limited availability of expert surgeons. Consequently, the demand for auto methods for personalized surgical mentorship stands as a critical, yet challenging issue. Despite their effectiveness in providing simple answers, these methods generally have difficulties when faced with challenging answer-location scenarios.
@@
Building upon the foundational groundwork of elementary Visual Question Answering (VQA) models,  [1, 2, 3, 4, 5, 6] innovatively delineates a paradigm for the community in the study of Surgical Visual Question Localized-Answering utilizing deep learning methodologies. However, these methods are mainly subject to two common limitations. Firstly, these methods have difficulty establishing accurate long-range dependencies when dealing with complex scenarios, due to occlusion and distortion, which may lead to the ignorance of relevant visual details that are significant to surgical environments. Additionally, they often overlook the alignment of multiple modalities, resulting in a mismatch between language and visual grounding. Considering these issues above, we ask: Can we personalize a large vision-language model to automatically ground VQA in surgical scenarios in a simple and efficient manner?
@@
With the rise of Large Language Models (LLMs) [7], Large Vision-Language Models (LVLMs) have emerged as a new method for understanding the interplay between visual and textual information [8, 9, 10]. However, existing large models still struggle to grasp domain-specific semantic information in subdomains [11], making a surgical-focused LVLM a current necessity. To this end, we present Surgical-LVLM, a personalized surgical large vision-language model for highly specialized surgical scenarios. As shown in Fig. 1(c), our method efficiently customizes LVLM faced with occlusive medical samples. Specifically, we first obtain the describable visual-language tokens by a pre-trained vision transformer (ViT) and Qwen-VL [12]. Consequently, we leverage a cross-attention layer stack of specially designed Visual Perception LoRA (VP-LoRA) blocks to unleash attention’s feature caption power with one core technique. Specifically, we integrate Visual State Space (VSS) blocks at each Low-Rank Adaptation layer, enabling the effective propagation of global contextual information and allowing Qwen-VL to better adapt to specialized surgical scenarios.
@@
With the aforementioned designs, Surgical-LVLM exhibits favorable grounding performance for hard samples in a variety of complex scenes. Nevertheless, there might be occasional failure cases, specifically the misalignment problem that the language model may describe some correct, yet subimportant content. Such ambiguity challenges Surgical-LVLM’s grounding capacity in determining the appropriate object to localize as output, since both the important and sub-important parts can be regarded as correct answers. To alleviate this issue, we further propose an alignment variant of our approach. In detail, we project the tokens generated from Qwen-VL to latent space, and take a token interaction (TIT) module to enable text tokens to emphasize the important vision information. By such efficient projection-based training, Surgical-LVLM selectively focuses on the relevant features or words that are most important for the VQLA task. Our contributions to this work are as follows:
@@
Figure 1: The flowchart of the proposed framework. The framework components include a pre-trained ViT, tokenizer, the Qwen-VL model, projection module, and the CAT-ViL embedding module.
II Related Work
II-A Large Visual-Language Model
@@
With the significant progress of large language models such as GPT-4 [7] and its open-source alternatives like LLaMA [13] and Vicuna [14], researchers have been increasingly motivated to exploit their integration of powerful language abilities with vision [15]. Recently, the performance of the general large multimodal models has been continuously refreshing the expectations since the scaling up of the pretraining data [12, 16] and instruction-following data [16, 17]. However, most models can only excel in general domains and easily encounter challenges in specific scenarios, such as complex surgical scenarios. In addition, while much of the current work has a good performance of language responses based on images and user instructions, it is difficult to provide alignment to related image regions (visual grounding) simultaneously. It is noted that CogVLM-Grounding [18] and MiniGPT-v2 [19] have a limitation in the generation of concise captions during grounding tasks, attributable to its primary training on caption data from datasets such as Flickr30K [20]. Conversely, BuboGPT [21] preserves conversational competency utilizing an external grounding model for grounding purposes, although the language encoder in the grounding model constrains the chat capability. These methodologies face challenges in achieving proficiency concurrently in both chat and grounding domains.
II-B VQLA Models in Medical Domain
@@
Medical VQLA has emerged as a pivotal area in the medical domain, leveraging the integration of textual questions and images to support clinical diagnostics and treatment. However, the unique characteristics inherent in medical data necessitate specialized expertise for accurate interpretation, thereby introducing additional complexities in developing effective medical VQLA models [22]. Recent advancements in medical VQLA research have been focused on addressing these inherent challenges. Surgical-VQLA [1] incorporates a gated vision-language embedding approach to fuse multimodal features from visual and textual sources. CAT-ViL [6] is subsequently proposed as a co-attention gated embedding module, allowing for instructive interaction between text and visual embeddings. These pioneering endeavors have contributed to several specialized improvements for medical scenarios and achieved superior representation learning. However, due to parameter limitations and the lack of specialized data, these models can just answer fixed-format questions and are unable to analyze comprehensive information about surgical scenarios to meet clinical needs. Therefore, the introduction of large language models [23, 13, 24] into the field of surgical-VQLA is necessitated by their powerful capabilities in text generation and comprehension.
III Methodology
III-A Preliminaries
III-A1 Qwen-VL
@@
is a Large Vision-Language Mode designed to perceive and understand both texts and images [12]. The model architecture consists of Qwen-7B as the language model, a Vision Transformer (ViT) as the visual encoder, and a vision-language adapter to compress the image features. During training, the model is trained in a three-stage pipeline: pretraining with image-text pairs, multi-task pretraining with interleaved vision-language data, and supervised fine-tuning. This training process enables Qwen-VL to interact with users, perceive input images, and generate appropriate responses based on user intent.
III-A2 Surgical-VQLA
@@
indicates models for Visual Question Localized-Answering in robotic surgery [1], which uses a combination of VisualBERT [2] and a Gated Vision-Language Embedding (GVLE) technique for effective fusion of visual and text features. The GVLE is designed to build input patches for the Language Vision Transformer and enables localized answering by incorporating a detection head in parallel with the prediction head of the LViT. Besides, CAT-ViL [6] further introduces the Co-Attention Gated Vision-Language (CAT-ViL) embedding module, which fuses multimodal features more effectively.
III-B Surgical-LVLM
@@
We propose Surgical-LVLM to perform complex scene comprehension as well as logical reasoning while performing VQLA tasks in surgical scenarios. As shown in Fig. 1, the framework consists of finetuned Qwen-VL (ViT [25], tokenizer, QwenLM and VP-LoRA blocks), a projection module and the grounding module (CAT-ViL).
III-B1 Visual Perception LoRA
@@
Due to problems such as occlusion and distortion, traditional Transformers have difficulties in establishing accurate long-range dependencies when dealing with images in complex surgical scenarios. Therefore, we design Visual Perception LoRA, which enhances the visual perception capabilities of Qwen-VL by integrating Visual State Space (VSS) blocks [26] at each Low-Rank Adaptation (LoRA) layer [27] during finetuning.
@@
As shown in Fig. 1 (a), the input first passes through an initial normalization layer. Then, the output is divided into two parallel information branches: the first branch contains linear transformation and activation, while the second branch incorporates linear transformation, depthwise separable convolution, and subsequent activation, followed by selective feature extraction through the 2D-Selective-Scan (SS2D) module. The SS2D is able to capture different features by expanding them and conducting feature extraction [28]. Finally, the linear layer is utilized to integrate the merged features of the two branches, which are then combined via the residual connection to generate the VP-LoRA output following the LoRA layer. With $X_{in}$ and $X_{out}$ representing the input and output features, the equation of the VP-LoRA block is as follows:
@@
III-B2 Multimodal Alignment for Visual Grounding
@@
In the alignment projection module, the language response is first processed through a linear layer and then combined with a residual connection. The combined features will be passed through the Token InTeraction (TIT) module [29] to further refine and emphasize the important information in the features. There are two parallel flows in the TIT module. The first flow contains two linear layers with an Aggregation block between them. With $x_{i}$ and $y_{i}$ representing the input and output features, the equations of the Aggregation block are as follows:
$\boldsymbol{y}_{i}=\mathcal{F}_{inter}\left[\mathcal{A}(i,\mathbf{X}),\boldsymbol{x}_{i}\right]$ (2)
where the contextual features for each location $i$ are aggregated using $\mathcal{A}$ first, and then the query is computed via interaction $\mathcal{F}_{inter}$ to generate $\boldsymbol{y}_{i}$. $\mathcal{A}$ denotes the Aggregation operation. The second branch is just linear transformation. Finally, the merged features of these two flows will be output to the CAT-ViL model for text embedding. The TIT module selectively focuses on the relevant features or words that are most important for the VQLA task.
@@
III-C Instruction Tuning
In order to enhance the performance of LVLMs’ capability for surgical scenarios understanding and logical reasoning, it is essential to create a dataset with high quality. For this purpose, we have established a dataset comprising question-answer pairs based on the images and annotations from the EndoVis 2017/2018 VQLA datasets. These pairs have been formatted according to the Qwen-VL structure. Guided by the methodology [30], we have formulated questions that span a spectrum of complexity, ranging from easy conversations, detailed descriptions, and complex reasoning as illustrated in Figure 2. These instruction-tuning QA pairs have been synthesized utilizing the GPT-4, which is renowned for its exemplary linguistic proficiency. The detailed statistical information on this dataset can be found in Sec IV-A.
III-D Training Strategy
@@
Vision-Language Instruction Fine-tuning: In the first phase, we leverage the extensive dataset comprising surgical question-answer pairs alongside associated images to fine-tune the Qwen-VL model augmented with VP-LoRA. This stage exclusively modifies the parameters of adapter layers while keeping the original large language model layers frozen. Through instruction fine-tuning, the Qwen-VL model acquires comprehension of the semantic intricacies and linguistic structures specific to the surgical domain from both textual and visual encoding features. Consequently, the model can generate meaningful representations, adeptly capturing the nuances and intricacies inherent in surgical terminology and context, thus establishing a robust foundation for effective vision-language grounding.
Visual Grounding Alignment: In the second stage of training, the fusion of chat and grounding capabilities is enacted by integrating the Qwen-VL and visual grounding module. Leveraging the pre-trained Qwen-VL model, language features extracted from its outputs are employed alongside corresponding surgical images to comprehensively optimize the parameters of CAT-ViL embedding. This optimization is geared towards enhancing the coherence between language responses and specific visual regions in surgical images. To achieve this alignment, we leverage CAT-ViL’s techniques with the co-attention embedding fusion to effectively combine the visual and textual features. Through this process, CAT-ViL acquires the capacity to accurately ground textual descriptions to pertinent regions in the surgical images by learning the associations between Qwen-VL’s output features and visual representations.
@@
IV Experiments
IV-A Dataset
EndoVis Conversations Dataset is a novel dataset generated from EndoVis-18-VQLA and EndoVis-17-VQLA Datasets. We designed questions encompassing three distinct types leveraging GPT-4. Statistically, the training and test sets include 19,020 and 2,151 image-QA pairs, respectively. The image setup for the training and test sets is also referenced to the above two datasets.
EndoVis-18-VQLA Dataset comprises 14 robotic surgery videos obtained from the MICCAI Endoscopic Vision Challenge [31]. Within this dataset, VQLA annotations are incorporated, integrating the question-answer pairs derived from [32] alongside bounding box annotations sourced from [33]. These QA pairs encompass a wide spectrum of single-word responses that encapsulate information regarding organs, surgical instruments, and their interrelations. We adhere to the data partitioning strategy delineated in [3] for the training and test datasets.
EndoVis-17-VQLA Dataset is emanated from the MICCAI Endoscopic Vision Challenge 2017 [34]. Leveraging this dataset as an external validation benchmark, we employ it to demonstrate the generalization efficacy of our proposed model across diverse surgical scenarios.
IV-B Implementation Details
We utilize the mainstream GPT-4 Score in the comparison experiments against BLIP-2 [35], miniGPT-4 [36], LLaVA-1.5 [23] and LLaVA-Med [37] on the EndoVis Conversations Dataset. Besides, for the EndoVis-18-VQLA and EndoVis-17-VQLA Datasets, the baseline models for our quantitative experimental comparison refer to [6]. The evaluation metrics are accuracy, f-score, and mean intersection over union (mIoU) [38]. All experiments are implemented by Python PyTorch framework, and conducted on a server with NVIDIA A100 GPUs. The epoch, batch size, and learning rate are set to 20, 16, and $1\times 10^{-5}$, respectively.
TABLE I: Ablation study on EndoVis Conversation dataset. ‘VP-LoRA’ denotes ‘Visual Perception LoRA’ and ‘IF’ denotes ‘Instruction Fine-tuning’.

### v3

@@ Abstract
Recent advancements in surgical Visual Question Answering (VQA) and related region grounding have shown great promise for robotic and medical applications, addressing the critical need for automated methods in personalized surgical mentorship. However, existing models primarily provide simple structured answers and struggle with complex scenarios due to their limited capability in recognizing long-range dependencies and aligning multimodal information. In this paper, we introduce Surgical-LVLM, a novel personalized large vision-language model tailored for complex surgical scenarios. Utilizing the pre-trained large vision-language model and specialized Visual Perception LoRA (VP-LoRA) blocks, our model excels in understanding complex visual-language tasks within surgical contexts. In addressing the visual grounding task, we propose the Token-Interaction (TIT) module, which strengthens the interaction between the grounding module and the language responses of the Large Visual Language Model (LVLM) after projecting them into the latent space. We demonstrate the effectiveness of Surgical-LVLM on several benchmarks, including EndoVis-17-VQLA, EndoVis-18-VQLA, and a newly introduced EndoVis Conversations dataset, which sets new performance standards. Our work contributes to advancing the field of automated surgical mentorship by providing a context-aware solution.
1 Introduction
@@
The progress achieved in surgical vision question answering (VQA) (Seenivasan et al., 2022a) and related region grounding has attracted many researchers since these techniques demonstrate great potential for applications in robotics and healthcare. To integrate the above techniques and achieve them simultaneously, Bai et al. (2023d) pioneered the surgical visual question localized-answering (VQLA), which predicts answers with localized results through the input text and images. Current approaches in this field predominantly rely on the expertise of expert surgeons, often failing to acknowledge the challenges posed by real-world constraints, such as the limited availability of expert surgeons. Consequently, the demand for auto methods for personalized surgical mentorship stands as a critical, yet challenging issue. Despite their effectiveness in providing simple answers, these methods generally have difficulties when faced with challenging answer-location scenarios.
@@
Building upon the groundwork of elementary Visual Question Answering (VQA) models,  (Seenivasan et al., 2023; Yuan et al., 2023; Bai et al., 2023b; c) innovatively delineates a paradigm for the community in the study of Surgical Visual Question Localized-Answering utilizing deep learning methodologies. However, these methods are mainly subject to two common limitations. Firstly, these methods have difficulty establishing accurate long-range dependencies when dealing with complex scenarios, due to occlusion and distortion, which may lead to the ignorance of relevant visual details that are significant to surgical environments. Additionally, they often overlook the alignment of multiple modalities, resulting in a mismatch between language and visual grounding. Considering these issues above, we ask: Can we personalize a large vision-language model to automatically ground VQA in surgical scenarios in a simple and efficient manner?
Figure 1: The flowchart of the proposed framework. It includes a pre-trained ViT, tokenizer, the Qwen-VL model, projection module, and the CAT-ViL embedding module.
@@
With the rise of Large Language Models (LLMs) (Achiam et al., 2023), Large Vision-Language Models (LVLMs) have emerged as a new method for understanding the interplay between visual and textual information (Liu et al., 2024a; Zhang et al., 2023b; You et al., 2023). However, existing large models still struggle to grasp domain-specific semantic information in subdomains (Li et al., 2024b), making a surgical-focused LVLM a current necessity. To this end, we present Surgical-LVLM, a personalized surgical large vision-language model for highly specialized surgical scenarios. As shown in Fig. 1(c), our method efficiently customizes LVLM faced with occlusive medical samples. Specifically, we first obtain the describable visual-language tokens by a pre-trained vision transformer (ViT) and Qwen-VL (Bai et al., 2023a). Consequently, we utilize a cross-attention layer stack of specially designed Visual Perception LoRA (VP-LoRA) blocks to unleash attention’s feature caption power with one core technique. Specifically, we integrate Visual State Space (VSS) blocks at each Low-Rank Adaptation layer, enabling the effective propagation of global contextual information and allowing Qwen-VL to better adapt to specialized surgical scenarios.
@@
With the aforementioned designs, Surgical-LVLM exhibits favorable grounding performance for hard samples in a variety of complex scenes. Nevertheless, there might be occasional failure cases, specifically the misalignment problem that the language model may describe some correct, yet subimportant content. Such ambiguity challenges Surgical-LVLM’s grounding capacity in determining the appropriate object to localize as output since both the important and sub-important parts can be regarded as correct answers. To alleviate this issue, we further propose an alignment variant of our approach. In detail, we project the tokens generated from Qwen-VL to latent space and take a token interaction (TIT) module to enable text tokens to emphasize the important vision information. By such efficient projection-based training, Surgical-LVLM selectively focuses on the relevant features or words that are most important for the VQLA task. Our contributions are as follows:
@@
2 Related Work
2.1 Large Visual-Language Model
@@
With the significant progress of LLMs such as GPT-4 (Achiam et al., 2023) and its open-source alternatives like LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023), researchers have been increasingly motivated to exploit their integration of powerful language abilities with vision (Zhang et al., 2023a). Recently, the performance of the general large multimodal models has been continuously refreshing the expectations since the scaling up of the pretraining data (Bai et al., 2023a; Dai et al., 2024) and instruction-following data (Dai et al., 2024; Li et al., 2023a; Wang et al., 2024). However, most models can only excel in general domains and easily encounter challenges in specific scenarios, such as complex surgical scenarios. In addition, while much of the current work has a good performance of language responses based on images and user instructions, it is difficult to provide alignment to related image regions (visual grounding) simultaneously. It is noted that CogVLM-Grounding (Wang et al., 2023) and MiniGPT-v2 (Chen et al., 2023) have a limitation in the generation of concise captions during grounding tasks, attributable to its primary training on caption data from datasets such as Flickr30K (Plummer et al., 2015). Conversely, BuboGPT (Zhao et al., 2023) preserves conversational competency utilizing an external grounding model for grounding purposes, although the language encoder in the grounding model constrains the chat capability. These methods face challenges in achieving proficiency in both chat and grounding domains.
2.2 VQLA Models in Medical Domain
@@
Medical VQLA has emerged as a pivotal area in the medical domain, using the integration of textual questions and images to support clinical diagnostics and treatment. However, the unique characteristics inherent in medical data necessitate specialized expertise for accurate interpretation, thereby introducing additional complexities in developing effective medical VQLA models (Liu et al., 2023b; Bai et al., 2025). Recent advancements in medical VQLA research have been focused on addressing these inherent challenges. Surgical-VQLA (Bai et al., 2023d) incorporates a gated vision-language embedding approach to fuse multimodal features from visual and textual sources. CAT-ViL (Bai et al., 2023b) is subsequently proposed as a co-attention gated embedding module, allowing for instructive interaction between text and visual embeddings. These pioneering endeavors have contributed to several specialized improvements for medical scenarios and achieved superior representation learning. However, due to parameter limitations and the lack of specialized data, these models can just answer fixed-format questions and are unable to analyze comprehensive information about surgical scenarios to meet clinical needs. Therefore, the introduction of large language models (Liu et al., 2023a; Touvron et al., 2023; Zheng et al., 2024) into the field of surgical-VQLA is necessitated by their powerful capabilities in text generation and comprehension.
3 Methodology
3.1 Preliminaries
3.1.1 Qwen-VL
@@
Qwen-VL is a Large Vision-Language Mode designed to perceive and understand both texts and images (Bai et al., 2023a). The model architecture consists of Qwen-7B as the language model, a Vision Transformer (ViT) as the visual encoder, and a vision-language adapter to compress the image features. During training, the model is trained in a three-stage pipeline: pretraining with image-text pairs, multi-task pretraining with interleaved vision-language data, and supervised fine-tuning. This training process enables Qwen-VL to interact with users, perceive input images, and generate appropriate responses based on user intent.
3.1.2 Surgical-VQLA
@@
Surgical-VQLA indicates models for Visual Question Localized-Answering in robotic surgery (Bai et al., 2023d), which uses a combination of VisualBERT (Li et al., 2019) and a Gated Vision-Language Embedding (GVLE) technique for effective fusion of visual and text features. The GVLE is designed to build input patches for the Language Vision Transformer and enables localized answering by incorporating a detection head parallel to the prediction head of the LViT. Besides, CAT-ViL (Bai et al., 2023b) further introduces the Co-Attention Gated Vision-Language (CAT-ViL) embedding module, which fuses multimodal features more effectively.
3.2 Surgical-LVLM
@@
We propose Surgical-LVLM to perform complex scene comprehension as well as logical reasoning while performing VQLA tasks in surgical scenarios. As shown in Fig. 1, the framework consists of finetuned Qwen-VL (ViT (Dosovitskiy et al., 2020), tokenizer, QwenLM and VP-LoRA blocks), a projection module and the grounding module (CAT-ViL).
3.2.1 Visual Perception LoRA
@@
Due to problems such as occlusion and distortion, traditional Transformers have difficulties in establishing accurate long-range dependencies when dealing with images in complex surgical scenarios. Therefore, we design Visual Perception LoRA, which enhances the visual perception capabilities of Qwen-VL by integrating Visual State Space (VSS) blocks (Liu et al., 2024c) at each Low-Rank Adaptation (LoRA) layer (Hu et al., 2021) during finetuning.
@@
As shown in Fig. 1 (a), the input first passes through an initial normalization layer. Then, the output is divided into two parallel information branches: the first branch contains linear transformation and activation, while the second branch incorporates linear transformation, depthwise separable convolution, and subsequent activation, followed by selective feature extraction through the 2D-Selective-Scan (SS2D) module. The SS2D is able to capture different features by expanding them and conducting feature extraction (Ruan & Xiang, 2024). Finally, the linear layer is utilized to integrate the merged features of the two branches, which are then combined via the residual connection to generate the VP-LoRA output following the LoRA layer. With $X_{in}$ and $X_{out}$ representing the input and output features, the equation of the VP-LoRA block is as follows:
@@
3.2.2 Multimodal Alignment for Visual Grounding
@@
In the alignment projection module, the language response is first processed through a linear layer and then combined with a residual connection. The combined features will be passed through the Token InTeraction (TIT) module (Yang et al., 2022) to further refine and emphasize the important information in the features. There are two parallel flows in the TIT module. The first flow contains two linear layers with an Aggregation block between them. With $x_{i}$ and $y_{i}$ representing the input and output features, the equations of the Aggregation block are as follows:
$\bm{y}_{i}=\mathcal{F}_{inter}\left[\mathcal{A}(i,\mathbf{X}),\bm{x}_{i}\right]$ (2)
where the contextual features for each location $i$ are aggregated using $\mathcal{A}$ first, and then the query is computed via interaction $\mathcal{F}_{inter}$ to generate $\bm{y}_{i}$. $\mathcal{A}$ denotes the Aggregation operation. The second branch is just linear transformation. Finally, the merged features of these two flows will be output to the CAT-ViL model for text embedding. The TIT module selectively focuses on the relevant features or words that are most important for the VQLA task.
@@
3.3 Instruction Tuning
In order to enhance the performance of LVLMs’ capability for surgical scenarios understanding and logical reasoning, it is essential to create a dataset with high quality. For this purpose, we have established a dataset comprising question-answer pairs based on the images and annotations from the EndoVis 2017/2018 VQLA datasets. These pairs have been formatted according to the Qwen-VL structure. Guided by the methodology (Liu et al., 2024b), we have formulated questions that span a spectrum of complexity, ranging from easy conversations, detailed descriptions, and complex reasoning as illustrated in Figure 2. These instruction-tuning QA pairs have been synthesized utilizing the GPT-4, which is renowned for its exemplary linguistic proficiency. The detailed statistical information on this dataset can be found in Sec 4.1.
3.4 Training Strategy
@@
Vision-Language Instruction Fine-tuning: In the first phase, we utilize the extensive dataset comprising surgical question-answer pairs alongside associated images to fine-tune the Qwen-VL model augmented with VP-LoRA. This stage exclusively modifies the parameters of adapter layers while keeping the original large language model layers frozen. Through instruction fine-tuning, the Qwen-VL model acquires comprehension of the semantic intricacies and linguistic structures specific to the surgical domain from both textual and visual encoding features. Thus, the model can generate meaningful representations, adeptly capturing the nuances and intricacies inherent in surgical terminology and context, thus establishing a robust foundation for effective vision-language grounding.
Visual Grounding Alignment: In the second stage of training, the fusion of chat and grounding capabilities is enacted by integrating the Qwen-VL and visual grounding module. Utilizing the pre-trained Qwen-VL model, language features extracted from its outputs are employed alongside corresponding surgical images to comprehensively optimize the parameters of CAT-ViL embedding. This optimization is geared towards enhancing the coherence between language responses and specific visual regions in surgical images. To achieve this alignment, we use the co-attention embedding fusion to effectively combine the multi-modal features. Through this process, CAT-ViL acquires the capacity to accurately ground textual descriptions to pertinent regions in the surgical images by learning the associations between Qwen-VL’s output features and visual representations.
@@
4 Experiments
4.1 Dataset
EndoVis Conversations Dataset is a novel dataset generated from EndoVis-18-VQLA and EndoVis-17-VQLA Datasets. We designed questions encompassing three distinct types using GPT-4. Statistically, the training and test sets include 19,020 and 2,151 image-QA pairs, respectively. The image setup for the training and test sets is also referenced to the above two datasets.
EndoVis-18-VQLA Dataset comprises 14 robotic surgery videos obtained from the MICCAI Endoscopic Vision Challenge (Allan et al., 2020). Within this dataset, VQLA annotations are incorporated, integrating the question-answer pairs derived from (Seenivasan et al., 2022b) alongside bounding box annotations sourced from (Islam et al., 2020). These QA pairs encompass a wide spectrum of single-word responses that encapsulate information regarding organs, surgical instruments, and their interrelations. We adhere to the data partitioning strategy delineated in (Seenivasan et al., 2022a) for the training and test datasets.
EndoVis-17-VQLA Dataset is emanated from the MICCAI Endoscopic Vision Challenge 2017 (Allan et al., 2019). Utilizing this dataset as an external validation benchmark, we employ it to demonstrate the generalization efficacy of our proposed model across diverse surgical scenarios.
4.2 Implementation Details
We utilize the mainstream GPT-4 Score in the comparison experiments against BLIP-2 (Li et al., 2023b), miniGPT-4 (Zhu et al., 2023), LLaVA-1.5 (Liu et al., 2023a) and LLaVA-Med (Li et al., 2024a) on the EndoVis Conversations Dataset. Besides, for the EndoVis-18-VQLA and EndoVis-17-VQLA Datasets, the baseline models for our quantitative experimental comparison refer to (Bai et al., 2023b). The evaluation metrics are accuracy, f-score, and mean intersection over union (mIoU) (Rezatofighi et al., 2019). All experiments are implemented by Python PyTorch framework, and conducted on a server with NVIDIA A100 GPUs. The epoch, batch size, and learning rate are set to 20, 16, and $1\times 10^{-5}$, respectively.
Table 1: Ablation study on EndoVis Conversation dataset. ‘VP-LoRA’ denotes ‘Visual Perception LoRA’ and ‘IF’ denotes ‘Instruction Fine-tuning’.